# Fairness is Not Silence: Unmasking Vacuous Neutrality in Small Language Models

### v1

@@ 4.1 How competent are Tiny and Small Language Models in reasoning under social bias scenarios?
Figure 2: F1 scores for (a) Tiny LMs (the first two heatmaps in blue) and (b) Small LMs (the last two heatmaps in green) across social bias categories. Rows correspond to bias categories and columns to instruction-tuned models. The heatmaps reflect F1 scores under Ambiguous context (first and third) and Disambiguated context (second and fourth). Darker shades reflect higher F1 scores and better task performance, while lighter shades indicate weaker competence.
Ideally, a fair and competent model should maintain or improve accuracy when constrained to provide unbiased responses. As demonstrated in Appendix Figure 6, the magnitude and direction of accuracy shifts under fairness constraints reveal how strongly a model depends on bias-aligned reasoning. LLaMA models show notable gains when forced to be fair, especially in stereotype-prone categories. For example, LLaMA3.2-1B improves by +11.0% on Disability Status and +13.5% on Physical Appearance. Even LLaMA3.2-3B shows gains (+8.3% on Religion, +8.8% on Gender Identity, +12.2% on Physical Appearance), suggesting that biased reasoning originally reduced task performance. Gemma models follow a similar, though milder, trend: Gemma3-1B improves on Religion (+2.1%), SES (+3.6%), and Disability Status (+8.3%), while Gemma3-4B shows gains on Nationality (+9.6%), Gender Identity (+5.6%), and Religion (+5.8%), but drops slightly on Disability Status (–2.4%), indicating possible trade-offs. In contrast, Qwen models exhibit virtually no change across categories, with values near zero. This suggests neutrality but also reflects prior observations that Qwen models may lack nuanced reasoning, relying neither on nor responding to demographic cues. The Phi family maintains strong performance with minimal reliance on biased patterns. Phi-3.5-mini improves in Religion (+9.0%) and Physical Appearance (+8.7%), while Phi-4-mini remains stable, with minor drops (e.g., –1.9% in Sexual Orientation), suggesting a well-balanced integration of fairness and competence.
Takeaways. Qwen family consistently yields near-zero bias scores in both ambiguous and disambiguous contexts, indicating a conservative generation strategy. While this results in favorable fairness metrics, it may come at the cost of specificity and task competence. In contrast, Phi models achieve both low bias and strong performance on reasoning tasks, even under ambiguity, demonstrating that fairness and utility can coexist when models are both informed and well-aligned.
5 How are model predictions distributed across answer options (A/B/C) and bias types (target, non-target and unknown)?
Figure 3: (Left) Target vs Non-Target Ratio: This plot shows the ratio of predictions favoring the biased (target) versus unbiased (non-target) outcome across social categories for small models. Values above 1.0 indicate stronger attribution to bias, while below 1.0 suggests bias denial. (Right) Proportion of “unknown”: this plot visualizes the ratio of model prediction "Unknown" responses to Ground Truth "Unknown" across social bias categories. High values indicate that a model refrains from committing to either biased or unbiased outcomes. A ratio near 1.0 indicates that the model correctly flags ambiguous cases as unresolvable.
By analyzing prediction distributions across answer choices and bias categories, we find that fairness arises from responsibly handling demographic cues rather than adopting a stance of vacuous neutrality, where models superficially appear unbiased by merely avoiding sensitive attributes. We investigate this using two key metrics: (1) the ratio of target versus non-target predictions across demographic categories, and (2) the proportion of predictions labeled as “unknown.” Specifically, we measure how closely each model’s frequency of “unknown” responses matches the ground truth frequency for each social bias category.
From Figure 3 (left), Phi (blue) and Gemma (purple) consistently maintain balanced ratios near 1.0 across most categories, signifying fairness. Conversely, Qwen (red) and LLaMA (green) show strong deviations. Qwen amplifies bias in Disability Status (154:1) and Physical Appearance (4:1), whereas LLaMA demonstrates strong bias denial, with ratios as low as 0.2–0.3 in these categories.
The right side of Figure 3 shows the ratio of "Unknown" predictions compared to the ground truth uncertainty. Phi exhibits highly calibrated uncertainty handling (0.8-0.9). For instance, in Race × Gender and Sexual Orientation, Phi achieves ratios close to 0.9, aligning well with the expected uncertainty rate. LLaMA is consistently overconfident, rarely using the "Unknown" option (0.1–0.2). Qwen demonstrates moderate caution (0.5–0.7), while Gemma scores lower (0.2–0.4). Notably, no model exceeds a ratio of 1.0, indicating general overconfidence rather than excessive uncertainty.
These findings are further substantiated by an analysis of the predicted label distributions (A/B/C) across social bias categories, accompanied by interpretation to clarify the observed trends, as detailed in Appendix B, Figures 8 through 11. Figure 9 reveals a systematic avoidance of Label A by the LLaMA3.2 models, a behavior consistent across model sizes. In contrast, Figure 8 highlights a persistent positional bias in Qwen variants, which tend to over-select Label A irrespective of model scale. For more balanced categories such as Gender Identity, Race and Ethnicity, and Religion, both Phi and Gemma families demonstrate greater fairness. Notably, Gemma3-4B exhibits stronger alignment with ground truth distributions compared to its smaller counterpart, as shown in Figure 10. Lastly, Figure 11 illustrates that Phi models employ a more balanced and context-sensitive label distribution, reflecting both a nuanced understanding of bias-sensitive contexts and consistent fairness across social dimensions.
Takeaways. LLaMA3.2-3B and Qwen2.5-3B often fail to interpret bias-related uncertainty effectively. In contrast, Phi-4-mini handles ambiguity well, maintains balanced responses, and reduces misclassification in sensitive cases, making it more suitable for fairness-critical applications.
6 What is the impact of model compression on the utility (F1 score) and fairness (bias score) of SLMs?
Model compression can sometimes enhance performance Lin et al. (2024b). Our previous analysis showed that uncompressed LLaMA models, particularly LLaMA3.2-3B, underperform compared to their peers in both competence and fairness. This raises a critical question: Does compression solely reduce model size, or can it also serve as an implicit regularizer that improves generalization and fairness in SLMs?
After applying AWQ 4-bit quantization Lin et al. (2024a), all models showed substantial size reductions, enabling efficient deployment on memory-constrained systems. Detailed compression statistics are presented in Appendix Table 3. To assess the impact of compression on bias-reasoning task, we calculate two key metrics: Relative F1 Change and Change in Bias. The formulas and detailed explanations of these metrics are provided in the Appendix C. These metrics help determine whether compression introduces trade-offs between utility and fairness or can simultaneously enhance both.
From Figure 4 (left) in the Appendix, we observe that Qwen2.5-3B exhibits uniformly light-colored cells in the heatmap, indicating that quantization had minimal effect. Its F1 performance remains stable across both ambiguous and disambiguated contexts. In contrast, Phi-4-Mini experiences substantial degradation across most bias categories, with deep red cells indicating F1 drops of 70–80% in both contexts. For example, F1 decreases by 67% in Age and by 72–80% in Disability Status, as well as significant losses in Gender Identity, Nationality, Race × Gender, and Race × SES. This indicates that compression severely affected Phi’s reasoning in bias-sensitive tasks. LLaMA3.2-3B shows a mixed outcome. In ambiguous contexts, it improves significantly after compression, with F1 increases of +118% on Age, +113% on Nationality, and up to +162% on Race × SES. However, in disambiguated contexts, it experiences mild declines, typically 3–12%, such as a 5.8% drop in Religion and 12% in Race × SES.
Similarly, as shown in Figure 4 (right) in the Appendix, we observe that Qwen2.5-3B shows minimal change post-compression, with most bias score differences within $\pm$0.5, indicating that quantization preserved its original fairness characteristics. In contrast, Phi-4-Mini exhibits significant and category-dependent shifts in bias. Compression altered its fairness properties both positively and negatively, with notable reductions in categories like SES, Religion, Race × Gender, and Physical Appearance under disambiguated contexts. However, it also caused sharp increases, most prominently in Disability Status, where bias rose by about 7.2%. LLaMA3.2-3B shows generally favorable changes, with decreased bias scores after compression, particularly in Physical Appearance and Religion categories where the original model displayed high bias. Although slight increases in bias are observed in specific instances, the overall trend indicates improved fairness, suggesting that compression can promote a more balanced model. These findings highlight that when compression techniques are appropriately aligned with the model’s architecture and training dynamics, they may preserve or even enhance fairness. Conversely, misalignment, as exemplified by the Phi models, can exacerbate bias, particularly in sensitive categories such as Disability Status.
As shown in Appendix Figure 7, the scatter plot illustrates how different models respond to compression across various bias categories and context types. The quadrant definitions clearly depict the trade-offs between utility and fairness introduced by compression. This visualization supports our RQ4 findings, showing that models like LLaMA3.2-3B generally benefit from compression in both utility and fairness, while Phi-4-Mini exhibits inconsistent behavior, sometimes reducing bias, but in other cases amplifying it.
Takeaways. The impact of compression on utility and fairness is inherently complex and varies significantly by model. While Phi-4-Mini suffers substantial performance loss and shows inconsistent fairness changes, improving in some areas but worsening in others, LLaMA3.2-3B not only preserves utility in ambiguous settings but also demonstrates a notable reduction in bias.
@@ 7 Conclusion
This work reveals that competence and fairness can coexist. While some models like Qwen2.5 appear neutral due to random or vacuous responses, demonstrating that fairness by silence is not a viable strategy. In contrast, the Phi family achieve high F1 scores ($\geq$ 90%) while remaining almost bias-free, showcasing the feasibility of lightweight, ethical NLP for edge deployments. In contrast, LLaMA3.2 models exhibit strong task performance but also pronounced stereotyping, which 4-bit AWQ quantization partially mitigates by reducing bias without sacrificing performance. These results underscore the importance of balanced evaluation, as high fairness scores may at times indicate model underperformance rather than genuine unbiased behavior. Considering both utility and fairness, our findings guide the development of efficient, capable, and socially responsible edge-ready language models.
@@ 8 Limitations
Our study is subject to several limitations that warrant consideration and present opportunities for future work. First, we restrict our analysis to open-source SLMs in the 0.5B–5B parameter range. Consequently, our conclusions about bias–capacity trade-offs are limited to this intermediate model scale and may not generalize outside this range, including proprietary models such as GPT-4 OpenAI et al. (2024). Second, our evaluation is limited to the BBQ dataset Parrish et al. (2022), which is well-designed for analyzing bias under context ambiguity but restricted to U.S.-centric social categories and a question-answering (QA) format. Extending this analysis to more diverse cultural contexts, additional languages, and broader downstream tasks such as summarization, dialogue, or retrieval, would enhance the generalizability of our findings. Finally, we consider only AWQ quantization as our compression method. Other techniques including structured/unstructured pruning and knowledge distillation may exhibit different effects on fairness and utility. As such, our findings should not be interpreted as representative of all compression strategies.
@@ 9 Ethical Considerations
Small language models (SLMs) enable fair, low-cost NLP on edge devices, increasing access and privacy. These models support on-device personalization and low-latency inference without cloud reliance. In effect, they can democratize advanced language technology (e.g. in healthcare or education) for resource-constrained or privacy-sensitive settings. As many of these models are produced using compression techniques, such methods can either obscure or amplify underlying biases. Moreover, a model that emits many neutral or “no-answer” responses may misleadingly appear fair (a phenomenon we call “vacuous neutrality”) while actually avoiding sensitive content. Such behavior yields representational harm: any systematic errors correlated with social identity (race, gender, disability, etc.) can reinforce stereotypes or exclude minorities. These considerations underline that true fairness requires examining both model competence and bias, silence or refusal alone is not an ethically adequate solution.
Appendix A BBQ Dataset and Evaluation Metrics
@@
The Bias Benchmark for Question Answering (BBQ) dataset Parrish et al. (2022) is a comprehensive benchmark designed to assess representational biases in language models. The BBQ dataset is licensed for non-commercial research use. All evaluated models are publicly available under open-source licenses (e.g., Apache 2.0, MIT) via HuggingFace. It comprises 58,492 unique question instances, each presented in both ambiguous and disambiguated formats. The dataset covers nine key demographic dimensions and two intersectional dimensions to facilitate a deeper examination of compound biases. Each question presents three answer choices: one that reflects a stereotypical bias (Target), one that challenges the stereotype (Non-Target), and an “Unknown” choice that reflects appropriate uncertainty. To evaluate model behavior, the original authors propose four metrics: accuracy on ambiguous questions (where the correct response is ideally “Unknown”), accuracy on disambiguated questions (where the model is expected to select the contextually appropriate answer), and two bias scores quantifying stereotypical tendencies under both ambiguous and disambiguated conditions. In this paper, we adopt the F1 score in place of accuracy to evaluate the model utility.
In disambiguated settings, the bias score $s_{\text{DIS}}$ quantifies the extent to which a model’s predictions align with social stereotypes. It is defined as follows:
where $n_{\text{biased-outputs}}$ is the number of predictions that conform to stereotypical expectations (e.g., choosing the Target in a negatively phrased question or the Non-Target in a positively phrased one), and $n_{\text{non-UNKNOWN-outputs}}$ denotes the number of predictions excluding those labeled as UNKNOWN. The value of $s_{\text{DIS}}$ ranges from $-100$ (fully anti-stereotypical) to $+100$ (fully stereotypical), with 00 indicating neutrality. For ambiguous contexts, the bias score $s_{\text{AMB}}$ incorporates both the degree of bias and the uncertainity of the model. It is computed as follows:
$s_{\text{AMB}}=(1-\text{accuracy})\cdot s_{\text{DIS}}$ (4)
Here, accuracy refers to the proportion of predictions where the model correctly chooses UNKNOWN in ambiguous scenarios. As a result, $s_{\text{AMB}}$ also falls within the range $[-100,+100]$, where values near zero indicate low bias or high uncertainty.

### v2

@@ 4.3 Evaluation Metrics
Figure 3: Heatmaps show F1 scores for (a) Tiny LMs (blue) and (b) Small LMs (green) under Ambiguous and Disambiguated contexts. Rows represent social bias categories and columns represent SLMs. Darker shades indicate higher F1 Score and stronger task performance; lighter shades denote weaker competence.
Utility Dimension For StereoSet and CrowS-Pairs, we evaluate utility using the Language Modeling Score (LMS) nadeem2020stereosetmeasuringstereotypicalbias, defined as the percentage of instances where the model favors a meaningful (stereotypical or anti-stereotypical) association over an unrelated one. An ideal model attains an LMS of 100. For BBQ, we measure task performance using the F1 score, computed separately for ambiguous and disambiguated contexts.
Ambiguity Handling Dimension The third dimension in the framework evaluates whether a model can abstain when appropriate (predicting Unknown) while still making specific predictions when sufficient context is provided. For StereoSet and CrowS-Pairs, ambiguity handling cannot be directly quantified, since ground truth labels only distinguish between stereo and anti-stereo completions and do not include explicit Unknown cases. For BBQ, we quantify ambiguity handling with two measures:
Target-to-NonTarget Ratio (TNR): the proportion of target predictions relative to non-target predictions, computed across the entire dataset in both ambiguous and disambiguated contexts (Eq. (3)).
Unknown Ratio (UR): the fraction of instances where the model predicts Unknown in ambiguous contexts, compared against the number of true Unknown instances (Eq. (3)). Together, these measures indicate whether a model balances caution with specificity, offering insight into its robustness under uncertainty.
$\text{TNR}=\frac{n_{\text{target}}}{n_{\text{nontarget}}},\qquad\text{UR}=\frac{n_{\text{predicted-UNK}}}{n_{\text{gold-UNK}}}$ (3)
Positional Bias Dimension The final dimension tests whether models favor certain answer positions {A, B, C} or stereotypical categories (stereo, anti-stereo, unknown). Such skews suggest reliance on heuristics rather than reasoning and can distort fairness and competence. We measure this using normalized Kullback–Leibler (KL) divergence between model predictions and a reference distribution. For BBQ, divergence is computed against the empirical ground truth distribution across positions. For StereoSet and CrowS-Pairs, where no distributional ground truth is provided, we can use a uniform reference distribution assuming equal probability across positions. We compute the normalized KL divergence, ranging from 0 to 1, with higher values indicating closer alignment to the reference distribution:
$\text{Norm-}D_{\text{KL}}(P\parallel Q)=1-\frac{\sum_{i}P(i)\log\frac{P(i)}{Q(i)}}{\log|C|}$ (4)
where $P(i)$ is the predicted probability for position $i$, $Q(i)$ is the ground truth or uniform distribution, and $|C|$ is the number of classes.
5 Experiments and Results
We present our experiments and results primarily for the BBQ benchmark, which natively supports all four dimensions of the VaNeu, including ambiguity handling. This makes BBQ the most comprehensive dataset for our analysis. Results on StereoSet and CrowS-Pairs, which focus on bias and utility, are discussed in more detail in the Appendix  B and  C respectively.
Bias Dimension
The first stage of our evaluation focuses on bias, asking whether models display systematic stereotypical preferences across demographic categories. Figure 2 reports bias scores across social categories in the BBQ dataset. Overall, most SLMs appear nearly unbiased, with all nine models registering within a narrow range of approximately ±15%. This indicates that none of the evaluated models exhibit extreme stereotypical alignment or strongly anti-stereotypical behavior. When grouped by family, distinct patterns emerge. The Qwen models consistently cluster near zero, reflecting a stable neutrality across contexts. The Phi family also maintains balanced bias levels, showing no systematic preference for stereotypical or anti-stereotypical completions. By comparison, the LLaMA and Gemma families display more variability across categories, occasionally reinforcing stereotypes but still remaining within the low-bias threshold. Stage 1 establishes a baseline where all nine models demonstrate low bias and meet responsible deployment standards, making them viable for Stage 2.
Figure 4: (Left) Target/Non-target Ratio (TNR) by category for SLMs; values >1.0>1.0 indicate a stronger tendency to predict target (stereotypical) over non-target, while values <1.0<1.0 indicate bias denial. (Middle) Unknown Ratio (UR): values 1.0 indicates that the model correctly flags ambiguous cases as unresolvable. (Right) Stage 4 positional bias measured as normalized KL divergence $({\text{Norm-}}D_{KL})$; higher is better and closer to the reference distribution. The dashed line marks the ground-truth baseline at $1.0$.
Utility Dimension
Stage 2 evaluates competence to carry out the QA task. Figure 3 shows that utility scores diverge much more sharply across families than bias alone. The LLaMA and Gemma models occupy a middle ground, their larger variants show strong gains under disambiguation, but tiny ones remain uneven and sometimes fall near random guessing. For example, LLaMA3.2-3B scores below 9% F1 on ambiguous Age and Nationality but exceeds 80% once demographic cues are explicit.
By contrast, the Phi family demonstrates that fairness and competence can align. Phi-3.5-Mini achieves over 90% F1 in ambiguous contexts, while Phi-4-Mini consistently surpasses 95% in disambiguated cases. This combination of robustness under ambiguity and strength with explicit cues makes the Phi series stand out as the most reliable across contexts, though both variants still show residual weakness on Physical Appearance. Finally, the Qwen family performs poorly, achieving only about 16% F1 in ambiguous contexts and marginally higher in disambiguated ones. Despite exhibiting near-zero bias in Stage 1, these results under both contexts show that the Qwen models underperform in this stage. This pattern exemplifies vacuous neutrality, models that appear unbiased by bias metrics but fail to deliver competent predictions. Based on Stage 2, Utility dimension, the models that remain viable for the next stage are: LLaMA3.2-3B, Gemma3-4B, Phi-3.5-Mini, and Phi-4-Mini.
Ambiguity Handling
The third stage assesses how models manage ambiguous inputs using two complementary metrics. In the left and center panels of the Figure 4 presents the target-to-nontarget ratio (TNR) and the unknown ratio (UR). Together they capture how well each model balances caution and specificity under uncertainty.
The Gemma3-4B model performs well in maintaining a balanced TNR, correctly distinguishing between target and non-target options, but fails to align with the ground-truth unknown ratio. This suggests that while Gemma3-4B can make confident predictions, it tends to overcommit even when ambiguity warrants abstention. The LLaMA3.2-3B model shows mixed behavior: in categories such as Age, Disability Status, and Physical Appearance, it tends to produce more anti-stereotypical responses, whereas in Gender Identity, Religion, and Race × Gender, it skews toward stereotypical outputs. This inconsistency indicates that LLaMA’s handling of ambiguity is highly category-dependent.
By contrast, the Phi family demonstrates strong robustness. Phi-4-Mini maintains a balanced target-to-nontarget ratio across most categories (except minor deviations in Disability Status and Religion) and aligns closely with the ground-truth unknown distribution, except for Age and Disability Status. This reflects an ability to abstain when necessary without compromising task competence. Phi-3.5-Mini exhibits similar and even stronger stability, though with slightly greater variability across categories. Based on Stage 3, Phi-3.5-Mini and Phi-4-Mini maintain balanced caution and specificity, advancing to the final stage.
Positional Bias
The final stage evaluates whether models favors certain answer positions rather than uniform distribution. Such tendencies indicate reliance on positional heuristics instead of genuine reasoning. Figure 4 (right) shows this behavior using $\text{Norm-}D_{KL}$}, and comparing model prediction distributions with ground truth baselines.
Both Phi models achieve values close to 1.0 across all social categories, indicating strong alignment with ground truth distributions and minimal positional skew. Phi-3.5-Mini shows slightly lower scores in categories such as Gender Identity and Physical Appearance, while Phi-4-Mini maintains near-perfect consistency. These results suggest that both models distribute attention appropriately across answer positions relying on content rather than positional or categorical shortcuts. Their near-ground-truth alignment reinforces that fairness and competence can coexist even in nuanced reasoning scenarios. Based on Stage 4, both Phi models exhibit minimal positional bias and maintain strong instruction following behavior.
6 Discussion
VaNeu Framework:
To address RQ1, we evaluate SLMs (0.5B–5B) across the four dimensions of the VaNeu. The staged analysis shows that models appearing fair may fail under tests of competence, uncertainty reasoning, or positional stability, highlighting the need for multidimensional fairness evaluation. In the Bias dimension, all nine models lie within ±15%, indicating minimal stereotyping. However, Stage 2 (Utility) reveals that low bias does not ensure competence, as many tiny models perform near chance, showing fairness alone has limited practical value.
If deployment were based only on Stages 1 and 2, we would risk releasing biased or unstable models. As discussed in Appendix A.2, Qwen2.5-3B initially appears deployable after Stage 1 but exhibits an extremely high TNR (153.86) in Disability Status Category and consistently low Norm-DKL (< 0.10) across categories, indicating overcommitment to a single option and a lack of meaningful differentiation. LLaMA3.2-3B performs in the utility under disambiguated contexts but fails in Stage 3, showing poor UR calibration and strong positional preference in Stage 4. Similarly, Gemma3-4B achieves high task utility in disambuigated contexts yet struggles with ambiguity handling. However, its Stage 4 answer distribution aligns more closely with the ground truth, suggesting that apparent neutrality stems from balanced outputs rather than genuine reasoning. We further tested the effect of task-specific fine-tuning (Appendix D); it improved disambiguated performance but reduced reasoning under ambiguity. Stages 3 and 4 refine the analysis by assessing specificity and distributional balance, revealing that fairness and utility must be interpreted jointly, as models prone to vacuous neutrality may appear reliable without genuine reasoning. We discussed the stage 3 and stage 4 results for Small LMs (2B-4B) in the Appendix  A.2
Fairness Behavior:
In view of RQ2, Physical Appearance consistently stands out as the most bias-sensitive category across the nine models. Gemma3-1B exhibit pronounced stereotypical alignment, with bias scores of +12.2% in ambiguous and +14.4% in disambiguated contexts. Latent cultural associations formed during pretraining often surface when models encounter references to non-normative traits (e.g., height, weight, etc.). SLMs demonstrate a 10–15% decline in utility and ambiguity handling for this category, indicating that entrenched stereotypes can directly impair task competence and contextual reasoning. To assess how model competence shifts under unbiased constraints in disambiguated contexts, we use the Bias Non-Alignment metric (Appendix A.1) to quantify the impact of stereotype alignment on task performance. Physical Appearance category shows consistent competence gains across multiple SLMs. In both the Age and Disability Status categories, bias behavior varies noticeably with model scale. Tiny variants tend to reinforce stereotypes, whereas their larger ones exhibit mildly anti-stereotypical nature, suggesting that increased model scale, often accompanied by more extensive instruction tuning, may introduce partial ethical calibration. However, this improvement in fairness does not translate to overall competence and reliable ambiguity handling: even in disambiguated contexts, SLMs continue to struggle with utility, reflecting difficulty in reasoning about socially sensitive attributes.
Meanwhile, categories such as SES, Gender Identity, and Nationality show moderate yet consistent bias patterns, largely stable across contexts and model sizes. Conversely, the Race-related categories and Sexual Orientation maintain consistently low bias even after disambiguation, while exhibiting strong utility and ambiguity handling-indicating balanced data representation and robust fairness alignment.
@@ 7 Conclusion
In this work, we presented the VaNeu Framework, a staged evaluation paradigm for assessing fairness and reliability in SLMs. By analyzing nine models across four families and multiple social bias categories, we demonstrated that low bias alone does not guarantee competence, robustness, or fair reasoning under ambiguity. Our findings reveal that SLMs often exhibit vacuous neutrality, appearing unbiased while lacking genuine understanding, highlighting the need for multidimensional evaluation before deployment. This framework provides a principled pathway for identifying such weaknesses and promoting responsible use of SLMs in socially sensitive contexts. As future work, we aim to mathematically formalize the concept of Vacuous Neutrality and develop a composite metric that consolidates the four evaluation dimensions into a single score, enabling standardized assessment of model bias and deployment suitability.
@@ 8 Limitations
Our study is subject to several limitations that warrant consideration and highlight avenues for future research. First, we focus exclusively on open-source SLMs within the 0.5B–5B parameter range. Consequently, our observations on bias–capacity trade-offs are limited to this intermediate scale and may not extend to larger or proprietary models such as GPT-4 (OpenAI et al., 2024). Second, our evaluation relies on the BBQ dataset (Parrish et al., 2022), which effectively probes bias under contextual ambiguity but remains confined to U.S.-centric social categories and a question-answering format. Extending the framework to multicultural contexts, different languages, additional architectures and broader downstream tasks such as summarization, dialogue, or retrieval would enhance generalizability. Finally, the notion of Vacuous Neutrality remains qualitatively defined; future work aims to formalize it mathematically and integrate the four stages into a unified composite metric.
@@ 9 Ethical Considerations
Small Language Models (SLMs) enable low-cost NLP on edge devices, enhancing access and privacy. By supporting on-device personalization and low-latency inference without cloud dependence, they help democratize advanced language technologies particularly in healthcare, education, and other resource-constrained or privacy-sensitive domains. However, because many SLMs rely on model compression techniques, such methods can either obscure or amplify underlying biases. Moreover, a model’s responses may appear fair along a single dimension while actually avoiding genuine reasoning, particularly in ambiguous situations. This vacuous neutrality behavior can lead to representational harm, as systematic errors correlated with social identities (e.g., race, gender, or disability) may reinforce stereotypes or marginalize groups. These considerations underscore that true fairness requires assessing beyond single dimension.
Appendix A BBQ Dataset
@@
The Bias Benchmark for Question Answering (BBQ) dataset parrish-etal-2022-bbq is a comprehensive benchmark designed to assess representational biases in language models. The BBQ dataset is licensed for non-commercial research use. All evaluated models are publicly available under open-source licenses (e.g., Apache 2.0, MIT) via HuggingFace. It comprises 58,492 unique question instances, each presented in both ambiguous and disambiguated formats. The dataset covers nine key demographic dimensions and two intersectional dimensions to facilitate a deeper examination of compound biases. Each question presents three answer choices: one that reflects a stereotypical bias (Target), one that challenges the stereotype (Non-Target), and an “Unknown” choice that reflects appropriate uncertainty. To evaluate model behavior, the original authors propose four metrics: accuracy on ambiguous questions (where the correct response is ideally “Unknown”), accuracy on disambiguated questions (where the model is expected to select the contextually appropriate answer), and two bias scores quantifying stereotypical tendencies under both ambiguous, $s_{\text{AMB}}$ and disambiguated conditions, $s_{\text{DIS}}$. In this paper, we adopt the F1 score in place of accuracy to evaluate the utility of the model. Both bias scores falls within the range [100,+100][-100,+100], where values near zero indicate low bias or neutral.
A.1 Bias Non-Alignment